OpenStudy (anonymous):

HELP ! What is (-m)^-3n if m=2 and n=-24 3 –3 4 -4

OpenStudy (anonymous):

\[(-m)^{-3}n=\frac{n}{(-m)^3}\]

OpenStudy (anonymous):

replace m by 2, n by -24 get \[\frac{-24}{(-2)^3}=\frac{-24}{-8}=3\]